Hypothesis Testing

Please type the answer by computer. (must answer and step degrees of freedom)

State SSW, SSB, the degrees of freedom, the F-statistics, and the test result.Three types of medium sized cars assembled in New Zealand have been test driven by a motoring magazine and compared on a vari
Answer #1

H0: Null Hypothesis: \mu 1=\mu 2=\mu 3

HA: Alternative Hypothesis: \mu 1\neq \mu 2\neq \mu 3 (At least one mean is different from the other 2 means)

From the given Table, the following Table is calculated:

Brand A Brand B Brand C Total
N 5 5 5 15
\sum X 40 43 49.5 132.5
Mean 40/5 =8 43/5 = 8.6 49.5/5=9.9 132.5/15=8.833
\sum X^{2} 320.64 372.06 491.41 1184.11
Std. Dev. 0.4 0.7517 0.5831 0.989

From the above Table, ANOVA Table is calculated as follows:

Source of Variation Sum of Squares Degrees of Freedom Mean Sum of Squares F
Between Treatments SSB=9.4333 \nu 1= 2 9.4333/2=4.7167 F - statistic = 4.7167/0.355=13.2864
Within treatments SSW=4.26 \nu 2 = 12 4.26/12=0.355
Total 13.6933 14

F - statistic = 4.7167/0.355=13.2864

Degrees of freedom for numerator = \nu 1 = 2

Degrees of freedom for denominator = \nu 2 = 12

By Technology, P - value = 0.00009

Since P - value = 0.0009 is less than \alpha = 0.05, the difference is significant. Reject null hypothesis.

Conclusion:

The data support the claim that there is a difference in the mean fuel consumption for the three makes of car.

Answers to questions asked:
(a) SSW=4.26

(b) SSB=9.4333

(c)

Degrees of freedom for numerator = \nu 1 = 2

Degrees of freedom for denominator = \nu 2 = 12

(d)

F - statistic =13.2864

(e) Test result:

The difference is significant. Reject null hypothesis.

Conclusion:

The data support the claim that there is a difference in the mean fuel consumption for the three makes of car.